A brother and sister try to communicate with a string tied between two tin cans

If the string is 9.5 m long, has a mass of 25g and is pulled taut with a tension of 7.0N , how much time does it take for a wave to travel from one end of the string to the other?
Answer #1

linear mass density of the string, mue = m/L

= 0.025/9.5

= 0.00263 kg/m

we know, wave speed on the string, v = sqrt(T/mue)

= sqrt(7/0.00263)

= 51.6 m/s

time taken for the wave to travel from one end to other end,

t = L/v

= 9.5/51.6

= 0.184 s <<<<<<----------Answer